### v1

# EmoGist: Efficient In-Context Learning for Visual Emotion Understanding

Abstract
In this paper, we introduce EmoGist, a training-free, in-context learning method for performing visual emotion classification with LVLMs. The key intuition of our approach is that context-dependent definition of emotion labels could allow more accurate predictions of emotions, as the ways in which emotions manifest within images are highly context dependent and nuanced. EmoGist pre-generates multiple explanations of emotion labels, by analyzing the clusters of example images belonging to each category. At test time, we retrieve a version of explanation based on embedding similarity, and feed it to a fast VLM for classification. Through our experiments, we show that EmoGist allows up to 13 points improvement in micro F1 scores with the multi-label Memotion dataset, and up to 8 points in macro F1 in the multi-class FI dataset.
EmoGist: Efficient In-Context Learning
for Visual Emotion Understanding
Ronald Seoh  and Dan Goldwasser Purdue University {bseoh, dgoldwas}@purdue.edu
1 Introduction
Automated classification of visual emotion (Ekman, 1993; Lang et al., 1999; Mikels et al., 2005) is an extremely challenging problem, as the ways in which emotions are embedded within images are inherently nuanced. Hence, even large-vision language models (LVLMs) that are extensively trained for reasoning over visual inputs struggle in detecting these emotions (Bhattacharyya and Wang, 2025), as their training may not necessarily involve the ability to understand such nuanced patterns.
In this paper, we introduce EmoGist, a training-free, in-context learning method for performing visual emotion classification with LVLMs. The key intuition of our approach is that the real meaning of different emotion labels could be dependent on the image’s context. For example, we could intuitively imagine that the way the emotion of ‘excitement’ for sporting events could be significantly different from the ‘excitement’ of the academics for an upcoming conference. Hence, guiding LVLMs with such context dependent definition of emotion labels could allow the models to better focus on the nuanced patterns of the image.
EmoGist automatically pre-generates nuanced, context-specific descriptions of emotion labels, by analyzing the clusters of example images belonging to each label. At test time, we retrieve a version of description based on embedding similarity, and feed it to a fast LVLM for final classification. Through our experiments, we show that EmoGist allows up to 13 points improvement in micro F1 scores with the multi-label emotion classification, and up to 8 points improvement in macro F1 for the multi-class case. We also demonstrate that EmoGist could achieve improvements in smaller VLMs with 2 billion parameters.
2 Related Work
While visual emotions has been extensively studied in the many related fields including computer vision and psychology (Mikels et al., 2005; Machajdik and Hanbury, 2010; Peng et al., 2015; You et al., 2016; Yang et al., 2023), we are only starting to see the efforts to exploit large vision-language models for automated understanding of visual emotions (Xie et al., 2024; Etesam et al., 2024; Xenos et al., 2024; Bhattacharyya and Wang, 2025).
Visual in-context learning (ICL) has seen considerable amount of interest in recent literatures (Zhang et al., 2023; Zhou et al., 2024; Zhang et al., 2024), where many work have investigated effective strategies for choosing visual ICL examples given the test instance. However, we believe that our work is first to investigate ICL strategies with LVLMs in detail for performing visual emotion understanding.
3 EmoGist
We describe the major components of EmoGist, our in-context learning method with LVLMs for emotion classification. Instead of naively retrieving individual examples based on the visual similarity of the image, the key idea is to obtain an nuanced explanation of emotion labels, which could effectively serve as the decision boundary for the LVLM to make its predictions on.
Because the same emotion could manifest in many different ways for across different images, we develop a strategy where we exploit stronger LVLMs to pre-generate multiple descriptions of different emotion labels, by analyzing the clusters of example images belonging to each category. At test time, we retrieve the version of explanation based on the similarity of the test image to cluster centroids, and feed it to a fast LVLM for final prediction.
Embedding and storing the pool of emotion label examples
In order to generate multiple descriptions of emotion labels, we begin by embedding the pool of example images with an embedding model. We use the mmE5 model, the state-of-the-art multimodal embedding model by Chen et al. (2025). Then we store the embeddings into a hnswlib vector database (Malkov and Yashunin, 2018).
Clustering
After creating the vector database of example images, we run the $k$-means clustering algorithm (Lloyd, 1982) against the set of embeddings to get different clusters. Because we are interested in creating multiple versions of explanations for specific labels, clustering is done separately for each emotion label. We tune the hyperparameter $k$ by setting a portion of example images aside as a validation set, and tune by evaluating the end task performance on them. Due to the limited computational resources available, we only experiment with the $k$ values of 2, 4, and 6.
Generating explanations
With the clustering information, we provide a strong LVLM with images from each cluster, and prompt them to explain why the given images belong to the emotion label. For our experiments, we use the Qwen2.5-VL 72B model (Bai et al., 2025) for generating explanations. As it is not possible to provide the LVLM with all images from the cluster due to GPU memory limits and context length, we select 4 images from each cluster to create one version of label explanation.
Ensembles
As we only select a subset of images from each cluster for generating an explanation, it may be the case that the selected images and the explanations may not sufficiently match the test image. In order to mitigate this issue, we introduce a simple ensemble scheme, where we generate multiple verisions of explanations for each cluster, perform multiple predictions against a single example and take the majority vote. We generate multiple explanations for the cluster by ranking all images within the cluster by their distance to the centroid, and generate explanations for every top 4 images in the ranked list.
Selection of clusters at test time
We note that EmoGist addresses both multi-label and multi-class classification cases. For multi-label classification, we assume that all test instances are binarized, and perform predictions for each candidate label given the image. For each candidate label, we retrieve the closest cluster among the candidates with the corresponding label. For multi-class classification where the classes are exclusive to each other, we perform the classification only once by providing the list of possible classes to the model. We perform the search across the entire set of clusters regardless of their classes, and use the closest cluster to the test image in terms of its distance to the centroid.
4 Experiments
For our main experiments, we use two datasets: one is the Memotion 1.0 dataset (Sharma et al., 2020; Jin et al., 2024), a collection of meme images from social media where we perform a multi-label classification across 4 labels: sarcastic, humorous, offensive, and motivational. The second dataset is the FI dataset (You et al., 2016), a collection of everyday images across the internet tagged based on the Ekman model (Ekman, 1993) of 8 classes: amusement, anger, awe, contentment, disgust, excitement, fear, and sadness. [^1] Please refer to Appendix A for more detailed dataset statistics.
4.1 Baselines
In order to save computational resources and make our discussions more clear, we choose three different SOTA LVLMs of similar sizes for our experiments: Qwen2.5-VL 7B (Bai et al., 2025), Aya Vision 8B (Dash et al., 2025), and InternVL2.5 8B-MPO (Wang et al., 2024). To ensure that our findings are not tied to particular pretraining, these three VLMs were chosen to ensure that they do not share the same image encoder or backbone LLM. We additionally provide a subset of results for smaller LVLMs in Table 2.
We compare EmoGist with the following comparable in-context learning methods:
  • Zero-Shot: We simply prompt the LVLM with the test image and prediction prompt.
  • Global Exp: Instead of providing example images to come up with an explanation, we prompt a large LVLM for “global" explanation, where we ask the model to describe the common features of the images with the query emotion label, without providing any references.
  • $\text{ICL}_{\text{sim}}$: For the multi-class case, we retrieve 4 images from the pool of example images based on their embedding similarity, regardless of their labels. This is closest to $\text{EmoGist}_{\text{n}}$, in terms of the number of example images used.
  • $\text{ICL}_{\text{all}}$: We also test more typical case of performing ICL for multiclass classification, where we provide one image each for all classes. Note that the FI dataset have 8 classes, resulting in 8 example images to be provided to a LVLM.
We also test two variants of EmoGist:
  • $\text{EmoGist}_{\text{n}}$: This variant of EmoGist uses 4 images from the cluster for explanation generation.
  • $\text{EmoGist}_{\text{e}}$: This variant of EmoGist performs ensembling, where we generate 3 versions of label explanation, each of them using 4 images without any overlap between them.
4.2 Results
Method Model Qwen2.5-VL 7B Aya Vision 8B InternVL2.5 8B-MPO Precision Recall F1 Precision Recall F1 Precision Recall F1 Zero-Shot 46.900 49.973 39.829 50.873 47.777 44.500 46.691 50.102 42.590 Global Exp 30.711 29.757 23.112 25.423 20.351 15.187 32.866 36.007 30.990 -16.189 -20.216 -16.717 -25.450 -27.426 -29.313 -13.825 -14.095 -11.600 $\text{ICL}_{\text{sim}}$ 48.802 50.009 42.810 51.568 42.174 36.068 42.482 43.563 37.605 +1.902 +0.036 +2.981 +0.695 -5.603 -8.432 -4.209 -6.539 -4.985 $\text{ICL}_{\text{all}}$ 18.765 14.541 5.300 6.863 12.511 1.192 16.763 16.518 13.565 -28.135 -35.432 -34.529 -44.010 -35.266 -43.308 -29.928 -33.584 -29.025 $\text{EmoGist}_{\text{n}}$ 52.752 51.968 47.754 52.342 51.013 47.020 52.284 51.495 47.490 +5.852 +1.995 +7.925 +1.469 +3.236 +2.520 +5.593 +1.393 +4.900 $\text{EmoGist}_{\text{e}}$ 52.609 51.779 47.719 52.437 51.026 47.055 52.971 51.934 48.019 +5.709 +1.806 +7.890 +1.564 +3.249 +2.555 +6.280 +1.832 +5.429
(a) FI
Zero-Shot 77.343 48.488 59.607 75.848 63.776 69.290 72.882 63.320 67.766 Global Exp 77.465 37.650 50.672 73.818 71.249 72.511 74.711 62.693 68.176 +0.122 -10.838 -8.935 -2.030 +7.473 +3.221 +1.829 -0.627 +0.410 $\text{ICL}_{\text{}}$ 75.033 64.461 69.346 76.597 60.867 67.832 72.239 55.220 62.593 -2.310 +15.973 +9.739 +0.749 -2.909 -1.458 -0.643 -8.100 -5.173 $\text{EmoGist}_{\text{n}}$ 75.650 64.689 69.742 66.621 83.115 73.959 72.539 73.987 73.256 -1.693 +16.201 +10.135 -9.227 +19.339 +4.669 -0.343 +10.667 +5.490 $\text{EmoGist}_{\text{e}}$ 78.328 66.800 72.106 72.758 85.625 78.669 75.783 73.189 74.463 +0.985 +18.312 +12.499 -3.090 +21.849 +9.379 +2.901 +9.869 +6.697
(b) Memotion
Table 1: Results of our methods and baselines. We report macro scores for FI and micro scores for Memotion. Boldfaces indicate the best performance for each metric across all methods. Green and red numbers indicate the performance changes over the zero-shot baseline.
EmoGist achieves robust performance gains over all the baselines. In Table 1, we can see that both $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ achieve consistent improvements over all the baselines. For FI, $\text{EmoGist}_{\text{e}}$ gains 7.89 points in terms of macro F1 score over Zero-Shot, and 4.91 points over $\text{ICL}_{\text{sim}}$. The trend is largely similar for Memotion as well, where $\text{EmoGist}_{\text{e}}$ gains 12.50 points in terms of micro F1 score over Zero-Shot, and 2.76 points over ICL. It is interesting to note that Global Exp, which is essentially providing the strong LVLM’s general knowledge about emotion labels, is considerably worse than the Zero-Shot baseline. Therefore, we could see that having EmoGist’s localized, cluster-specific label explanation makes a substantial difference. Lastly, adding ensembling $\text{EmoGist}_{\text{e}}$ shows consistent improvements over $\text{EmoGist}_{\text{n}}$, with notable gains in precision.
Naively performing visual ICL could be detrimental. Another observation from Table 1 is that our ICL baselines, $\text{ICL}_{\text{sim}}$ and $\text{ICL}_{\text{all}}$, are either underperforming, or marginally better than the Zero-Shot baselines. In the most extreme case, $\text{ICL}_{\text{all}}$ for FI sees over 43 points drop in F1 score, way below the random guess. In addition, while ICL achieves considerable performance with Qwen2.5-VL 7B on Memotion, Aya Vision 8B and InternVL 2.5 8B-MPO failed to achieve comparable scores. As reasoning over multiple image inputs is still an area of active research in pretraining and post-training for LVLMs (Li et al., 2024), it is likely the case that not all publicily available LVLMs are equal in terms of leveraging ICL examples for classification.
Model Method FI Memotion Precision Recall F1 Precision Recall F1 SmolVLM2 2.2B Zero-Shot 39.660 30.743 20.712 73.148 67.598 70.264 ICL 47.575 48.536 41.825 75.000 11.808 20.404 +7.915 17.793 21.113 +1.852 -55.790 -49.860 $\text{EmoGist}_{\text{e}}$ 52.182 50.700 46.437 62.702 99.544 76.940 +12.522 +19.957 +25.725 -10.446 +31.946 +6.676 InternVL2.5 2B-MPO Zero-Shot 38.936 35.661 31.104 73.960 54.763 62.930 ICL 45.386 10.473 15.475 64.150 75.128 69.207 +6.450 -25.188 -15.629 -9.810 +20.365 +6.277 $\text{EmoGist}_{\text{e}}$ 50.987 42.970 41.670 70.612 60.582 65.213 +12.051 +7.309 +10.566 -3.348 +5.819 +2.283
Table 2: Results on small VLMs with 2B parameters.
Small VLMs could also become decent emotion reasoners with EmoGist. As many practical uses of visual emotion understanding often take place within resource-constrained systems with low latency requirements such as web applications or personal computing devices, even relatively small 7 billion models may be beyond typical computing budget under such scenarios. In Table 2, we test 2 small VLMs with the same number of 2 billion parameters, SmolVLM2 2.2B (Marafioti et al., 2025) and InternVL2.5 2B-MPO (Wang et al., 2024), to examine whether the performance benefits acheived by EmoGist still hold for these smaller models. We could see that $\text{EmoGist}_{\text{e}}$ achieve similar levels of performance gains over the Zero-Shot and ICL baselines, with SmolVLM2.2 achieving performances similar to the 7B models in Table 1. Given that EmoGist only requires storing cluster centroids and text descriptions at test time, we believe that EmoGist shows some interesting future directions for implementing visual emotion understanding into a wide variety of applications.
Model Method ArtPhoto Precision Recall F1 Qwen2.5-VL 7B Zero-Shot 52.594 42.306 41.178 ICL 43.332 33.165 33.745 EmoGist 46.601 38.555 39.229 Aya Vision 8B Zero-Shot 55.602 43.706 43.473 ICL 43.328 26.902 26.742 EmoGist 46.216 38.307 38.921 InternVL2.5 8B-MPO Zero-Shot 48.824 42.365 43.518 ICL 39.606 32.083 33.018 EmoGist 47.941 39.465 40.311
Table 3: Results on ArtPhoto with the emotion label explanations from FI.
Knowledge transfer across different domains with EmoGist. Lastly, we explore whether the knowledge about different emotion labels we acquire from example images could be used for predictions against the images from the domains different from example images. We use the ArtPhoto dataset, In Table 3, we can see that while EmoGist achieves slighty better scores over naive ICL, the overall performance is actually worse than the Zero-Shot baselines. Making EmoGist to capture more general knowledge of emotions across different subject domains and visual compositions could be an interesting direction for future research.
5 Conclusion and Future Work
In this paper, we introduced EmoGist, a training-free in-cotext learning method for visual emotion understanding with large vision-language models. We observe a significant amount of improvements over the zero-shot and naive ICL baselines across SOTA LVLMs of 2 and 7 billion parameters. In particular, we find that $\text{EmoGist}_{\text{e}}$, the variant of our method with a simple ensembling, achieves robust performance improvements with higher precision. In future work, we’d like to explore more deeply into the label explanations generated by the strong LVLMs and investigate various reasoning strategies for obtaining emotion label descriptions that are more transferrable across different domains.
Limitations
Due to limited computational resources available to the authors, we perform our experiments on a limited subset of publicly available large vision-language models with 2 and 7 billion parameters. While we anticipate that our findings would hold overall for other model sizes, we do not provide any direct evidence. We also note that we only tried one model each for the embedding model and the explanation generation LVLM, as stated in section 3.
Appendix A Dataset Information
For each dataset, we treat the training set as the pool of example images for explanation generation, use the validation set for hyperparameter tuning, and test our method and baselines against the test set.
A.1 Memotion
We use the split introduced by Jin et al. (2024) for our experiments. [^2] https://huggingface.co/datasets/Ahren09/MMSoc_Memotion. Dataset statistics are provided below:
Split Number of unique images sarcastic humorous offensive motivational Training 5593 4367 4259 3417 1972 Validation 699 538 539 437 253 Test 700 543 543 425 242
Table 4: Memotion dataset statistics.
A.2 FI
For the test set, we use the FI hard set introduced by (Bhattacharyya and Wang, 2025). As there is no separate training and validation set provided, we filter all the images includes in the FI hard set from the original FI set, and randomly split the set of remaining images into the training and validation set. Dataset statistics are provided below:
Split amusement anger awe contentment disgust excitement fear sadness Training 3608 842 2698 4309 1391 2592 852 2634 Validation 190 45 142 259 74 137 45 139 Test 1125 368 293 188 192 185 149 128
Table 5: FI dataset statistics.
Appendix B Hardwares and Softwares Used
To fully utilize all the GPU resources available to us, we ran all of test predictions using multiple NVIDIA V100, GTX TITAN Xp and GTX TITAN X GPUs.
We use the version 4.50.0 of HuggingFace Transformers library (Wolf et al., 2020), the version 0.7.3 of vLLM (Kwon et al., 2023) alongside PyTorch version 2.5.1 (Paszke et al., 2019).
For our label explanation generation, we used a Mac Studio hardware with M1 Ultra CPU and 128GB of RAM, using the version 0.25.0 of mlx (Hannun et al., 2023) and 0.1.23 of mlx-vlm.
Appendix C Prompts used for label description generation
These are the examples of images with the sentiment of "{s_label}". Based on these examples, what are the common features of images to be felt "{s_label}" by its viewers? Do not reference example images directly.
Appendix D Prompts used for classification
D.1 Multi-label
[…]
D.2 Multi-class
[…]

### v2

@@
 The key intuition of our approach is that context-dependent definition of emotion labels could allow more accurate predictions of emotions, as the ways in which emotions manifest within images are highly context dependent and nuanced.
-EmoGist pre-generates multiple explanations of emotion labels, by analyzing the clusters of example images belonging to each category.
-At test time, we retrieve a version of explanation based on embedding similarity, and feed it to a fast VLM for classification.
-Through our experiments, we show that EmoGist allows up to 13 points improvement in micro F1 scores with the multi-label Memotion dataset, and up to 8 points in macro F1 in the multi-class FI dataset.
+EmoGist pre-generates multiple descriptions of emotion labels, by analyzing the clusters of example images belonging to each label.
+At test time, we retrieve a version of description based on the cosine similarity of test image to cluster centroids, and feed it together with the test image to a fast LVLM for classification.
+Through our experiments, we show that EmoGist allows up to 12 points improvement in micro F1 scores with the multi-label Memotion dataset, and up to 8 points in macro F1 in the multi-class FI dataset.
 EmoGist: Efficient In-Context Learning
@@
 1 Introduction
+Figure 1: For visual emotion classification, naive in-context learning (ICL) struggles as providing multiple nuanced visual examples often lead LVLMs to make incorrect predictions.
+EmoGist guides LVLMs using pre-generated multiple descriptions of emotion labels obtained by analyzing the clusters of example images.
 Automated classification of visual emotion (Ekman, 1993; Lang et al., 1999; Mikels et al., 2005) is an extremely challenging problem, as the ways in which emotions are embedded within images are inherently nuanced.
-Hence, even large-vision language models (LVLMs) that are extensively trained for reasoning over visual inputs struggle in detecting these emotions (Bhattacharyya and Wang, 2025), as their training may not necessarily involve the ability to understand such nuanced patterns.
+Hence, even large vision-language models (LVLMs) that are extensively trained for reasoning over visual inputs struggle in detecting these emotions (Bhattacharyya and Wang, 2025), as their training may not necessarily involve the ability to understand such nuanced patterns.
 In this paper, we introduce EmoGist, a training-free, in-context learning method for performing visual emotion classification with LVLMs.
@@
 EmoGist automatically pre-generates nuanced, context-specific descriptions of emotion labels, by analyzing the clusters of example images belonging to each label.
-At test time, we retrieve a version of description based on embedding similarity, and feed it to a fast LVLM for final classification.
-Through our experiments, we show that EmoGist allows up to 13 points improvement in micro F1 scores with the multi-label emotion classification, and up to 8 points improvement in macro F1 for the multi-class case.
-We also demonstrate that EmoGist could achieve improvements in smaller VLMs with 2 billion parameters.
+At test time, we retrieve a version of description based on the cosine similarity of test image to cluster centroids, and feed it together with the test image to a fast LVLM for classification.
+Through our experiments, we show that EmoGist allows up to 12 points improvement in micro F1 scores with the multi-label emotion classification, and up to 8 points improvement in macro F1 for the multi-class case.
+We also demonstrate that EmoGist could achieve improvements in smaller LVLMs with 2 billion parameters.
+[^1] All the program codes used to produce results presented in this paper are available at https://tinyurl.com/emo-gist.
 2 Related Work
-While visual emotions has been extensively studied in the many related fields including computer vision and psychology (Mikels et al., 2005; Machajdik and Hanbury, 2010; Peng et al., 2015; You et al., 2016; Yang et al., 2023), we are only starting to see the efforts to exploit large vision-language models for automated understanding of visual emotions (Xie et al., 2024; Etesam et al., 2024; Xenos et al., 2024; Bhattacharyya and Wang, 2025).
-Visual in-context learning (ICL) has seen considerable amount of interest in recent literatures (Zhang et al., 2023; Zhou et al., 2024; Zhang et al., 2024), where many work have investigated effective strategies for choosing visual ICL examples given the test instance.
-However, we believe that our work is first to investigate ICL strategies with LVLMs in detail for performing visual emotion understanding.
+While visual emotions has been extensively studied in the many related fields including computer vision and psychology (Mikels et al., 2005; Machajdik and Hanbury, 2010; Peng et al., 2015; You et al., 2016; Yang et al., 2023), we are only starting to see the efforts to exploit large vision-language models for automated understanding of visual emotions (Xie et al., 2024; Etesam et al., 2024; Xenos et al., 2024; Lei et al., 2025; Bhattacharyya and Wang, 2025).
+Visual in-context learning (ICL) has seen considerable amount of interest in recent literature (Zhang et al., 2023; Zhou et al., 2024; Zhang et al., 2024), where many work have investigated effective strategies for choosing visual ICL examples given the test instance.
+However, we believe that our work is first to investigate ICL strategies with LVLMs in detail for evoked emotion classification.
 3 EmoGist
 We describe the major components of EmoGist, our in-context learning method with LVLMs for emotion classification.
-Instead of naively retrieving individual examples based on the visual similarity of the image, the key idea is to obtain an nuanced explanation of emotion labels, which could effectively serve as the decision boundary for the LVLM to make its predictions on.
-Because the same emotion could manifest in many different ways for across different images, we develop a strategy where we exploit stronger LVLMs to pre-generate multiple descriptions of different emotion labels, by analyzing the clusters of example images belonging to each category.
-At test time, we retrieve the version of explanation based on the similarity of the test image to cluster centroids, and feed it to a fast LVLM for final prediction.
+Instead of naively retrieving individual examples based on the visual similarity of the image, the key idea is to obtain an nuanced description of emotion labels, which could effectively serve as the decision boundary for the LVLM to make its predictions on.
+Because the same emotion could manifest in many different ways for across different images, we develop a strategy where we utilize stronger LVLMs to pre-generate multiple descriptions of different emotion labels, by analyzing the clusters of example images belonging to each category.
+Figure 2: Example label descriptions for clusters within the contentment class of the FI dataset, as generated by the Qwen2.5-VL 72B model.
 Embedding and storing the pool of emotion label examples
 In order to generate multiple descriptions of emotion labels, we begin by embedding the pool of example images with an embedding model.
-We use the mmE5 model, the state-of-the-art multimodal embedding model by Chen et al.
+We use the MM-E5 model, the state-of-the-art multimodal embedding model by Chen et al.
 (2025).
@@
 After creating the vector database of example images, we run the $k$-means clustering algorithm (Lloyd, 1982) against the set of embeddings to get different clusters.
-Because we are interested in creating multiple versions of explanations for specific labels, clustering is done separately for each emotion label.
+Because we are interested in creating multiple versions of descriptions for specific labels, clustering is done separately for each emotion label.
 We tune the hyperparameter $k$ by setting a portion of example images aside as a validation set, and tune by evaluating the end task performance on them.
-Due to the limited computational resources available, we only experiment with the $k$ values of 2, 4, and 6.
-Generating explanations
-With the clustering information, we provide a strong LVLM with images from each cluster, and prompt them to explain why the given images belong to the emotion label.
-For our experiments, we use the Qwen2.5-VL 72B model (Bai et al., 2025) for generating explanations.
-As it is not possible to provide the LVLM with all images from the cluster due to GPU memory limits and context length, we select 4 images from each cluster to create one version of label explanation.
-Ensembles
-As we only select a subset of images from each cluster for generating an explanation, it may be the case that the selected images and the explanations may not sufficiently match the test image.
-In order to mitigate this issue, we introduce a simple ensemble scheme, where we generate multiple verisions of explanations for each cluster, perform multiple predictions against a single example and take the majority vote.
-We generate multiple explanations for the cluster by ranking all images within the cluster by their distance to the centroid, and generate explanations for every top 4 images in the ranked list.
+Due to the limited computational resources available, we only experiment with the $k$ values of 2, 4, and 6.222Please see Appendix C for hyperparameter tuning procedures and sensitivity analysis.
+Generating label descriptions
+With the cluster information, we provide a strong LVLM with images from each cluster, and prompt them to explain why the given images belong to the emotion label.
+For our experiments, we use the Qwen2.5-VL 72B model (Bai et al., 2025) for generating descriptions.
+As it is not possible to provide the LVLM with all images from the cluster due to GPU memory limits and context length, we select 4 images from each cluster to create one version of label description.
+Figure 2 shows the prompt used and examples of generated descriptions.
 Selection of clusters at test time
@@
 For each candidate label, we retrieve the closest cluster among the candidates with the corresponding label.
-For multi-class classification where the classes are exclusive to each other, we perform the classification only once by providing the list of possible classes to the model.
+For multi-class classification where the classes are exclusive to each other, we perform the classification only once by providing the list of all candidate classes to the model.
 We perform the search across the entire set of clusters regardless of their classes, and use the closest cluster to the test image in terms of its distance to the centroid.
+Once the cluster and associated label description has been chosen, we prepend the description to the test image and classification prompt.
+[^3] Please see Appendix E for the full prompts used for classification.
+Ensembles
+As we only select a subset of images from each cluster for generating an description, it may be the case that the selected images and descriptions may not sufficiently match the test image.
+In order to mitigate this issue, we introduce a simple ensemble scheme, where we generate multiple versions of descriptions for each cluster, perform multiple predictions against a single example and take the majority vote.
+We generate multiple descriptions for the cluster by ranking all images within the cluster by their distance to the centroid, and generate descriptions for every top 4 images in the ranked list.
 4 Experiments
-For our main experiments, we use two datasets: one is the Memotion 1.0 dataset (Sharma et al., 2020; Jin et al., 2024), a collection of meme images from social media where we perform a multi-label classification across 4 labels: sarcastic, humorous, offensive, and motivational.
+Method Model Qwen2.5-VL 7B Aya Vision 8B InternVL2.5 8B-MPO Precision Recall F1 Precision Recall F1 Precision Recall F1 Zero-Shot 47.121 50.381 40.089 50.094 47.181 44.188 47.237 50.379 42.939 $\pm$0.065 $\pm$0.090 $\pm$0.065 $\pm$0.285 $\pm$0.188 $\pm$0.102 $\pm$0.297 $\pm$0.171 $\pm$0.223 Global Exp 30.503 30.156 23.480 26.416 20.782 15.405 32.242 35.224 30.583 $\pm$0.325 $\pm$0.439 $\pm$0.404 $\pm$1.670 $\pm$0.468 $\pm$0.322 $\pm$0.316 $\pm$0.452 $\pm$0.312 -16.619 -20.226 -16.609 -23.678 -26.398 -28.783 -14.996 -15.155 -12.355 $\text{ICL}_{\text{sim}}$ 48.656 49.898 42.733 48.649 42.444 35.257 42.613 43.658 37.653 $\pm$0.029 $\pm$0.022 $\pm$0.016 $\pm$0.592 $\pm$0.122 $\pm$0.185 $\pm$0.215 $\pm$0.171 $\pm$0.161 +1.534 -0.483 +2.643 -1.445 -4.736 -8.931 -4.624 -6.721 -5.286 $\text{ICL}_{\text{all}}$ 23.464 14.633 5.664 13.223 12.609 1.511 16.682 16.107 13.454 $\pm$0.948 $\pm$0.024 $\pm$0.081 $\pm$3.242 $\pm$0.042 $\pm$0.102 $\pm$0.517 $\pm$0.259 $\pm$0.298 -23.657 -35.749 -34.425 -36.871 -34.572 -42.677 -30.555 -34.272 -29.485 $\text{EmoGist}_{\text{n}}$ 52.944 52.163 48.497 52.579 51.389 47.906 52.592 51.767 48.094 $\pm$0.364 $\pm$0.265 $\pm$0.572 $\pm$0.381 $\pm$0.271 $\pm$0.565 $\pm$0.261 $\pm$0.220 $\pm$0.303 +5.822 +1.782 +8.408 +2.485 +4.208 +3.718 +5.354 +1.388 +5.155 $\text{EmoGist}_{\text{e}}$ 52.772 51.704 48.118 52.632 51.146 47.579 52.449 51.712 47.795 $\pm$0.377 $\pm$0.235 $\pm$0.552 $\pm$0.371 $\pm$0.316 $\pm$0.599 $\pm$0.158 $\pm$0.202 $\pm$0.167 +5.650 +1.323 +8.029 +2.538 +3.965 +3.391 +5.211 +1.333 +4.856
+(a) FI
+Zero-Shot 77.343 48.260 59.434 75.799 63.814 69.292 73.520 63.377 68.070 $\pm$0.013 $\pm$0.104 $\pm$0.080 $\pm$0.023 $\pm$0.065 $\pm$0.035 $\pm$0.195 $\pm$0.358 $\pm$0.254 Global Exp 77.134 35.206 48.159 73.430 68.159 70.665 74.885 62.160 67.928 $\pm$0.180 $\pm$2.182 $\pm$1.976 $\pm$0.702 $\pm$0.654 $\pm$0.116 $\pm$0.191 $\pm$0.360 $\pm$0.234 -0.209 -13.054 -11.276 -2.369 +4.345 +1.373 +1.365 -1.217 -0.142 $\text{ICL}_{\text{sim}}$ 75.061 64.642 69.462 76.721 60.915 67.910 72.601 55.476 62.890 $\pm$0.025 $\pm$0.107 $\pm$0.071 $\pm$0.043 $\pm$0.077 $\pm$0.062 $\pm$0.101 $\pm$0.432 $\pm$0.290 -2.282 +16.381 +10.028 +0.922 -2.900 -1.382 -0.918 -7.901 -5.181 $\text{EmoGist}_{\text{n}}$ 75.610 62.693 68.540 68.693 84.265 75.611 72.162 72.438 72.289 $\pm$0.183 $\pm$0.492 $\pm$0.237 $\pm$0.548 $\pm$1.958 $\pm$0.747 $\pm$0.204 $\pm$0.708 $\pm$0.338 -1.732 +14.432 +9.105 -7.107 +20.451 +6.319 -1.358 +9.061 +4.219 $\text{EmoGist}_{\text{e}}$ 78.682 65.374 71.411 70.263 87.165 77.795 75.898 74.596 75.240 $\pm$0.067 $\pm$0.315 $\pm$0.188 $\pm$0.314 $\pm$0.647 $\pm$0.171 $\pm$0.257 $\pm$0.173 $\pm$0.170 +1.339 +17.114 +11.977 -5.536 +23.350 +8.502 +2.379 +11.219 +7.170
+(b) Memotion
+Table 1: Results of our methods and baselines.
+We report macro scores for FI and micro scores for Memotion.
+All scores are averaged over six random seeds.
+We show standard errors as their confidence intervals.
+Boldfaces indicate the best performance for each metric across all methods for each model.
+Green and red numbers indicate the performance changes over the zero-shot baseline.
+We used two datasets: one is the Memotion 1.0 dataset (Sharma et al., 2020; Jin et al., 2024), a collection of meme images from social media where we perform a multi-label classification across 4 labels: sarcastic, humorous, offensive, and motivational.
 The second dataset is the FI dataset (You et al., 2016), a collection of everyday images across the internet tagged based on the Ekman model (Ekman, 1993) of 8 classes: amusement, anger, awe, contentment, disgust, excitement, fear, and sadness.
-[^1] Please refer to Appendix A for more detailed dataset statistics.
+[^4] Please see Appendix A for more detailed dataset statistics.
 4.1 Baselines
@@
 We compare EmoGist with the following comparable in-context learning methods:
-  • Zero-Shot: We simply prompt the LVLM with the test image and prediction prompt.
-  • Global Exp: Instead of providing example images to come up with an explanation, we prompt a large LVLM for “global" explanation, where we ask the model to describe the common features of the images with the query emotion label, without providing any references.
-  • $\text{ICL}_{\text{sim}}$: For the multi-class case, we retrieve 4 images from the pool of example images based on their embedding similarity, regardless of their labels.
-This is closest to $\text{EmoGist}_{\text{n}}$, in terms of the number of example images used.
-  • $\text{ICL}_{\text{all}}$: We also test more typical case of performing ICL for multiclass classification, where we provide one image each for all classes.
+  • Zero-Shot: We simply prompt a LVLM with the test image and prediction prompt.
+  • Global Exp: Instead of providing example images for a description, we prompt a large LVLM for “global" description, where we ask the model to describe the common features of the images with the candidate emotion label, without providing any references.
+  • $\text{ICL}_{\text{sim}}$: We retrieve 4 images from the pool of images based on cosine similarity, regardless of their labels.
+This is closest to $\text{EmoGist}_{\text{n}}$, in terms of the number of examples.
+  • $\text{ICL}_{\text{all}}$: We also test another case of performing ICL for multiclass classification, where we provide one image each for all classes.
 Note that the FI dataset have 8 classes, resulting in 8 example images to be provided to a LVLM.
 We also test two variants of EmoGist:
-  • $\text{EmoGist}_{\text{n}}$: This variant of EmoGist uses 4 images from the cluster for explanation generation.
-  • $\text{EmoGist}_{\text{e}}$: This variant of EmoGist performs ensembling, where we generate 3 versions of label explanation, each of them using 4 images without any overlap between them.
+  • $\text{EmoGist}_{\text{n}}$: This variant of EmoGist uses 4 images from the cluster for description generation.
+  • $\text{EmoGist}_{\text{e}}$: This variant of EmoGist performs ensembling, where we generate 3 versions of label description, each of them using 4 images without any overlap between them.
 4.2 Results
-Method Model Qwen2.5-VL 7B Aya Vision 8B InternVL2.5 8B-MPO Precision Recall F1 Precision Recall F1 Precision Recall F1 Zero-Shot 46.900 49.973 39.829 50.873 47.777 44.500 46.691 50.102 42.590 Global Exp 30.711 29.757 23.112 25.423 20.351 15.187 32.866 36.007 30.990 -16.189 -20.216 -16.717 -25.450 -27.426 -29.313 -13.825 -14.095 -11.600 $\text{ICL}_{\text{sim}}$ 48.802 50.009 42.810 51.568 42.174 36.068 42.482 43.563 37.605 +1.902 +0.036 +2.981 +0.695 -5.603 -8.432 -4.209 -6.539 -4.985 $\text{ICL}_{\text{all}}$ 18.765 14.541 5.300 6.863 12.511 1.192 16.763 16.518 13.565 -28.135 -35.432 -34.529 -44.010 -35.266 -43.308 -29.928 -33.584 -29.025 $\text{EmoGist}_{\text{n}}$ 52.752 51.968 47.754 52.342 51.013 47.020 52.284 51.495 47.490 +5.852 +1.995 +7.925 +1.469 +3.236 +2.520 +5.593 +1.393 +4.900 $\text{EmoGist}_{\text{e}}$ 52.609 51.779 47.719 52.437 51.026 47.055 52.971 51.934 48.019 +5.709 +1.806 +7.890 +1.564 +3.249 +2.555 +6.280 +1.832 +5.429
-(a) FI
-Zero-Shot 77.343 48.488 59.607 75.848 63.776 69.290 72.882 63.320 67.766 Global Exp 77.465 37.650 50.672 73.818 71.249 72.511 74.711 62.693 68.176 +0.122 -10.838 -8.935 -2.030 +7.473 +3.221 +1.829 -0.627 +0.410 $\text{ICL}_{\text{}}$ 75.033 64.461 69.346 76.597 60.867 67.832 72.239 55.220 62.593 -2.310 +15.973 +9.739 +0.749 -2.909 -1.458 -0.643 -8.100 -5.173 $\text{EmoGist}_{\text{n}}$ 75.650 64.689 69.742 66.621 83.115 73.959 72.539 73.987 73.256 -1.693 +16.201 +10.135 -9.227 +19.339 +4.669 -0.343 +10.667 +5.490 $\text{EmoGist}_{\text{e}}$ 78.328 66.800 72.106 72.758 85.625 78.669 75.783 73.189 74.463 +0.985 +18.312 +12.499 -3.090 +21.849 +9.379 +2.901 +9.869 +6.697
-(b) Memotion
-Table 1: Results of our methods and baselines.
-We report macro scores for FI and micro scores for Memotion.
-Boldfaces indicate the best performance for each metric across all methods.
-Green and red numbers indicate the performance changes over the zero-shot baseline.
 EmoGist achieves robust performance gains over all the baselines.
 In Table 1, we can see that both $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ achieve consistent improvements over all the baselines.
-For FI, $\text{EmoGist}_{\text{e}}$ gains 7.89 points in terms of macro F1 score over Zero-Shot, and 4.91 points over $\text{ICL}_{\text{sim}}$.
-The trend is largely similar for Memotion as well, where $\text{EmoGist}_{\text{e}}$ gains 12.50 points in terms of micro F1 score over Zero-Shot, and 2.76 points over ICL.
+For FI, $\text{EmoGist}_{\text{n}}$ gains 8.41 points in terms of macro F1 score over Zero-Shot, and 5.76 points over $\text{ICL}_{\text{sim}}$.
+The trend is largely similar for Memotion, where $\text{EmoGist}_{\text{e}}$ gains 11.98 points in terms of micro F1 score over Zero-Shot, and 1.95 points over $\text{ICL}_{\text{sim}}$.
 It is interesting to note that Global Exp, which is essentially providing the strong LVLM’s general knowledge about emotion labels, is considerably worse than the Zero-Shot baseline.
-Therefore, we could see that having EmoGist’s localized, cluster-specific label explanation makes a substantial difference.
-Lastly, adding ensembling $\text{EmoGist}_{\text{e}}$ shows consistent improvements over $\text{EmoGist}_{\text{n}}$, with notable gains in precision.
+Therefore, we could see that having EmoGist’s localized, cluster-specific label description makes a substantial difference.
+Lastly, adding ensembling $\text{EmoGist}_{\text{e}}$ shows consistent improvements over $\text{EmoGist}_{\text{n}}$, with notable gains in precision for Memotion.
 Naively performing visual ICL could be detrimental.
-Another observation from Table 1 is that our ICL baselines, $\text{ICL}_{\text{sim}}$ and $\text{ICL}_{\text{all}}$, are either underperforming, or marginally better than the Zero-Shot baselines.
-In the most extreme case, $\text{ICL}_{\text{all}}$ for FI sees over 43 points drop in F1 score, way below the random guess.
+Another observation from Table 1 is that our ICL baselines, $\text{ICL}_{\text{sim}}$ and $\text{ICL}_{\text{all}}$, are either under-performing, or marginally better than the Zero-Shot baselines.
+In the most extreme case, $\text{ICL}_{\text{all}}$ for FI sees over 42 points drop in F1 score, way below random guessing.
 In addition, while ICL achieves considerable performance with Qwen2.5-VL 7B on Memotion, Aya Vision 8B and InternVL 2.5 8B-MPO failed to achieve comparable scores.
-As reasoning over multiple image inputs is still an area of active research in pretraining and post-training for LVLMs (Li et al., 2024), it is likely the case that not all publicily available LVLMs are equal in terms of leveraging ICL examples for classification.
-Model Method FI Memotion Precision Recall F1 Precision Recall F1 SmolVLM2 2.2B Zero-Shot 39.660 30.743 20.712 73.148 67.598 70.264 ICL 47.575 48.536 41.825 75.000 11.808 20.404 +7.915 17.793 21.113 +1.852 -55.790 -49.860 $\text{EmoGist}_{\text{e}}$ 52.182 50.700 46.437 62.702 99.544 76.940 +12.522 +19.957 +25.725 -10.446 +31.946 +6.676 InternVL2.5 2B-MPO Zero-Shot 38.936 35.661 31.104 73.960 54.763 62.930 ICL 45.386 10.473 15.475 64.150 75.128 69.207 +6.450 -25.188 -15.629 -9.810 +20.365 +6.277 $\text{EmoGist}_{\text{e}}$ 50.987 42.970 41.670 70.612 60.582 65.213 +12.051 +7.309 +10.566 -3.348 +5.819 +2.283
-Table 2: Results on small VLMs with 2B parameters.
-Small VLMs could also become decent emotion reasoners with EmoGist.
+As reasoning over multiple image inputs is still an area of active research in pretraining and post-training for LVLMs (Li et al., 2024), it is likely the case that not all publicly available LVLMs are equal in terms of their ability to utilize ICL examples for classification.
+Small LVLMs could also become decent emotion reasoners with EmoGist.
 As many practical uses of visual emotion understanding often take place within resource-constrained systems with low latency requirements such as web applications or personal computing devices, even relatively small 7 billion models may be beyond typical computing budget under such scenarios.
-In Table 2, we test 2 small VLMs with the same number of 2 billion parameters, SmolVLM2 2.2B (Marafioti et al., 2025) and InternVL2.5 2B-MPO (Wang et al., 2024), to examine whether the performance benefits acheived by EmoGist still hold for these smaller models.
+In Table 2, we test 2 small LVLMs with the same number of 2 billion parameters, SmolVLM2 2.2B (Marafioti et al., 2025) and InternVL2.5 2B-MPO (Wang et al., 2024), to examine whether EmoGist performance benefits hold for these smaller models.
 We could see that $\text{EmoGist}_{\text{e}}$ achieve similar levels of performance gains over the Zero-Shot and ICL baselines, with SmolVLM2.2 achieving performances similar to the 7B models in Table 1.
 Given that EmoGist only requires storing cluster centroids and text descriptions at test time, we believe that EmoGist shows some interesting future directions for implementing visual emotion understanding into a wide variety of applications.
-Model Method ArtPhoto Precision Recall F1 Qwen2.5-VL 7B Zero-Shot 52.594 42.306 41.178 ICL 43.332 33.165 33.745 EmoGist 46.601 38.555 39.229 Aya Vision 8B Zero-Shot 55.602 43.706 43.473 ICL 43.328 26.902 26.742 EmoGist 46.216 38.307 38.921 InternVL2.5 8B-MPO Zero-Shot 48.824 42.365 43.518 ICL 39.606 32.083 33.018 EmoGist 47.941 39.465 40.311
-Table 3: Results on ArtPhoto with the emotion label explanations from FI.
+Model Method FI Memotion Precision Recall F1 Precision Recall F1 InternVL2.5 2B-MPO Zero-Shot 38.039 34.813 30.499 74.272 54.972 63.181 $\pm$0.285 $\pm$0.320 $\pm$0.318 $\pm$0.164 $\pm$0.129 $\pm$0.113 $\text{ICL}_{\text{sim}}$ 45.376 11.215 16.078 64.674 75.157 69.522 $\pm$0.587 $\pm$0.221 $\pm$0.309 $\pm$0.104 $\pm$0.090 $\pm$0.074 +7.337 -23.598 -14.421 -9.599 +20.184 +6.341 $\text{EmoGist}_{\text{e}}$ 50.706 41.421 37.442 70.873 60.658 65.353 $\pm$0.412 $\pm$0.442 $\pm$0.675 $\pm$0.142 $\pm$0.840 $\pm$0.504 +12.667 +6.608 +6.943 -3.400 +5.685 +2.173 SmolVLM2 2.2B Zero-Shot 39.567 30.762 20.703 73.124 67.617 70.263 $\pm$0.056 $\pm$0.019 $\pm$0.016 $\pm$0.011 $\pm$0.012 $\pm$0.010 $\text{ICL}_{\text{sim}}$ 47.478 48.463 41.798 74.896 12.113 20.852 $\pm$0.030 $\pm$0.029 $\pm$0.028 $\pm$0.097 $\pm$0.087 $\pm$0.130 +7.912 +17.700 +21.096 +1.773 -55.505 -49.411 $\text{EmoGist}_{\text{e}}$ 52.358 50.641 46.713 67.455 96.311 79.329 $\pm$0.542 $\pm$0.098 $\pm$0.445 $\pm$0.381 $\pm$0.540 $\pm$0.118 +12.791 +19.879 +26.010 -5.668 +28.694 +9.066
+Table 2: Results on small VLMs with 2B parameters.
 Knowledge transfer across different domains with EmoGist.
 Lastly, we explore whether the knowledge about different emotion labels we acquire from example images could be used for predictions against the images from the domains different from example images.
-We use the ArtPhoto dataset, In Table 3, we can see that while EmoGist achieves slighty better scores over naive ICL, the overall performance is actually worse than the Zero-Shot baselines.
-Making EmoGist to capture more general knowledge of emotions across different subject domains and visual compositions could be an interesting direction for future research.
+Using the label descriptions obtained from the example images of the FI dataset, we evaluate $\text{EmoGist}_{\text{e}}$ on the ArtPhoto dataset (Machajdik and Hanbury, 2010), a collection of artistically photographed images annotated with the same class labels as FI.
+In Table 3, we can see that while EmoGist achieves slightly better scores over naive ICL, the overall performance is actually worse than the Zero-Shot baselines.
+As EmoGist captures more context-specific, nuanced knowledge of emotion labels, there seems to be a significant semantic gap between the images from FI and ArtPhoto that most of the FI clusters do not adequately explain the test images from ArtPhoto.
+Making EmoGist to capture both general and context-specific knowledge of emotions across different subject domains and visual compositions could be an interesting direction for future research.
+Model Method ArtPhoto Precision Recall F1 Qwen2.5-VL 7B Zero-Shot 52.594 42.306 41.178 $\text{ICL}_{\text{sim}}$ 43.332 33.165 33.745 $\text{EmoGist}_{\text{e}}$ 46.535 38.555 39.227 Aya Vision 8B Zero-Shot 55.602 43.706 43.473 $\text{ICL}_{\text{sim}}$ 43.328 26.902 26.742 $\text{EmoGist}_{\text{e}}$ 46.216 38.307 38.921 InternVL2.5 8B-MPO Zero-Shot 48.824 42.365 43.518 $\text{ICL}_{\text{sim}}$ 39.606 32.083 33.018 $\text{EmoGist}_{\text{e}}$ 47.739 39.104 39.993
+Table 3: Results on ArtPhoto with the emotion label descriptions from FI.
 5 Conclusion and Future Work
@@
 We observe a significant amount of improvements over the zero-shot and naive ICL baselines across SOTA LVLMs of 2 and 7 billion parameters.
-In particular, we find that $\text{EmoGist}_{\text{e}}$, the variant of our method with a simple ensembling, achieves robust performance improvements with higher precision.
-In future work, we’d like to explore more deeply into the label explanations generated by the strong LVLMs and investigate various reasoning strategies for obtaining emotion label descriptions that are more transferrable across different domains.
+In particular, we find that $\text{EmoGist}_{\text{e}}$, the variant of our method with simple ensembling, achieves robust performance improvements with higher precision.
+In future work, we’d like to explore more deeply into the label descriptions generated by the strong LVLMs and investigate various reasoning strategies for obtaining emotion label descriptions that are more transferrable across different domains.
+Acknowledgments
+The work was supported by NSF CAREER award IIS2048001 and the DARPA CCU program. Contents do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
 Limitations
@@
 While we anticipate that our findings would hold overall for other model sizes, we do not provide any direct evidence.
-We also note that we only tried one model each for the embedding model and the explanation generation LVLM, as stated in section 3.
+We also note that we only tried one model each for the embedding model and the description generation LVLM, as stated in section 3.
 Appendix A Dataset Information
-For each dataset, we treat the training set as the pool of example images for explanation generation, use the validation set for hyperparameter tuning, and test our method and baselines against the test set.
+For each dataset, we treat the training set as the pool of example images for description generation, use the validation set for hyperparameter tuning, and test our method and baselines against the test set.
 A.1 Memotion
@@
 Table 5: FI dataset statistics.
-Appendix B Hardwares and Softwares Used
-To fully utilize all the GPU resources available to us, we ran all of test predictions using multiple NVIDIA V100, GTX TITAN Xp and GTX TITAN X GPUs.
+Appendix B Experimental Settings
+B.1 Hardwares and Softwares Used
+To fully utilize all the GPU resources available to us, we ran all of our validation and test predictions, and embedding generations using multiple NVIDIA V100, GTX TITAN Xp and GTX TITAN X GPUs.
 We use the version 4.50.0 of HuggingFace Transformers library (Wolf et al., 2020), the version 0.7.3 of vLLM (Kwon et al., 2023) alongside PyTorch version 2.5.1 (Paszke et al., 2019).
-For our label explanation generation, we used a Mac Studio hardware with M1 Ultra CPU and 128GB of RAM, using the version 0.25.0 of mlx (Hannun et al., 2023) and 0.1.23 of mlx-vlm.
-Appendix C Prompts used for label description generation
-These are the examples of images with the sentiment of "{s_label}".
-Based on these examples, what are the common features of images to be felt "{s_label}" by its viewers?
+For our label description generation, we used a Mac Studio hardware with M1 Ultra CPU and 128GB of RAM, using the version 0.25.0 of mlx (Hannun et al., 2023) and 0.1.23 of mlx-vlm.
+B.2 Random Seeds
+For the results shown in Table 1 and Table 2, we run each test for 6 different random seeds: 21, 42, 63, 84, 105, 126.
+Appendix C Sensitivity analysis and hyperparameter tuning for the number of clusters $k$
+We determine the optimal number of clusters for each dataset and test-time LVLM combination.
+Initially, we apply $k$-means clustering to both FI and Memotion datasets for $k$ values of 2, 4, and 6.
+Subsequently, we generate cluster-based label descriptions for each $k$ using the Qwen2.5-VL 72B model.
+We then classify the validation set based on these clusters and text descriptions, 3 times for each 7B/2B LVLM using the following random seeds: 21, 42, 63.
+The optimal $k$ is ultimately identified by identifying $k$ that achieved the most number of highest validation F1 scores across the three seeds.
+If there’s no winner, we ran the validation for additional seeds (84, 105, 126) until the clear winner is found.
+Please see Figure 9 and Figure 10 for the validation set results of $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ with 7B models on FI, Figure 11 and Figure 12 for the validation set results of $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ with 7B models on Memotion, and Figure 13 and Figure 14 for the validation set results of $\text{EmoGist}_{\text{e}}$ with 2B models on FI and Memotion.
+Appendix D Prompts used for label description generation
+**Images from the cluster go here**
+These are the examples of images with the sentiment of “{s_label}".
+Based on these examples, what are the common features of images to be felt “{s_label}" by its viewers?
 Do not reference example images directly.
-Appendix D Prompts used for classification
-D.1 Multi-label
-[…]
-D.2 Multi-class
-[…]
+Figure 3: Prompts used for label description generation.
+Appendix E Prompts used for zero-shot and EmoGist
+**Test image goes here**
+**For EmoGist, label description goes here**
+Question: Does this image match the sentiment label ‘{s_label}’? Answer:
+Answer with ‘Yes’ or ‘No’.
+Figure 4: Prompts used for multi-label zero-shot classification and EmoGist.
+**Test image goes here**
+**For EmoGist, label description goes here**
+Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
+Answer with the exact sentiment label as it appears in the list.
+Figure 5: Prompts used for multi-class zero-shot classification and EmoGist.
+Appendix F Prompts used for ICL baselines
+**ICL example images 1 to 4 goes here**
+**Test image goes here**
+## Image 1
+Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
+## Image 2
+Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
+## Image 3
+Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
+## Image 4
+Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
+## Image 5
+Question: Does this image matches the sentiment label ’{s_label}’? Answer:
+Answer with ‘Yes’ or ‘No’.
+Figure 6: Prompts used for multi-label $\text{ICL}_{\text{sim}}$.
+**ICL example images 1 to 4 goes here**
+**Test image goes here**
+## Image 1
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
+## Image 2
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
+## Image 3
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
+## Image 4
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
+## Image 5
+Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
+Answer with the exact sentiment label as it appears in the list.
+Figure 7: Prompts used for multi-class $\text{ICL}_{\text{sim}}$.
+**ICL example images 1 to 8 goes here**
+**Test image goes here**
+## Image 1
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: amusement
+## Image 2
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: anger
+## Image 3
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: awe
+## Image 4
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: contentment
+## Image 5
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: disgust
+## Image 6
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: excitement
+## Image 7
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: fear
+## Image 8
+Question: Which of the sentiment labels in the following list does this image belong to?
+List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: sadness
+## Image 9
+Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
+Answer with the exact sentiment label as it appears in the list.
+Figure 8: Prompts used for multi-class, $\text{ICL}_{\text{all}}$.
+Figure 9: Validation set results for $\text{EmoGist}_{\text{n}}$ with 7B models on FI.
+Figure 10: Validation set results for $\text{EmoGist}_{\text{e}}$ with 7B models on FI.
+Figure 11: Validation set results for $\text{EmoGist}_{\text{n}}$ with 7B models on Memotion.
+Figure 12: Validation set results for $\text{EmoGist}_{\text{e}}$ with 7B models on Memotion.
+Figure 13: Validation set results for $\text{EmoGist}_{\text{e}}$ with 2B models on FI.
+Figure 14: Validation set results for $\text{EmoGist}_{\text{e}}$ with 2B models on Memotion.